# Onsager reciprocal relations for charge and spin transport in periodically driven systems

### v1

@@ IV Floquet linear-response theory
and performing the similar procedures to those used to determine $[G^{r}_{a\sigma b\sigma^{\prime}}(\mathbf{k},\omega)]_{mn}$’s. Namely, the group velocities and Green’s functions appearing in Eqs. (89) and (90) are calculated with $\mathbf{A}_{\textrm{pump}}(-t)$ in the similar way to those with $\mathbf{A}_{\textrm{pump}}(t)$. For $[\bar{\epsilon}_{ab}(\mathbf{k})]_{mn}$ of Sr2RuO4 driven by BCPL, see Appendix C.
IV.3 General remarks about the applicability
@@
Then, our theory has wider applicability than the theories using a high-frequency expansion. This expansion has been often used to analyze many periodically driven systems Floquet-review1 ; Highw1 ; Light-AHE1 ; Mikami ; NA-BCPL . It may be sufficient for a periodically driven electron system if the light frequency is much larger than the bandwidth of the system. Meanwhile, our theory does not have such a restriction because the Floquet theory used here is free from the constraint on the frequency.
@@
The applicability of our theory is also wider than that of the theories in which the time-averaged off-diagonal dc conductivities are expressed in terms of the Berry curvature. In the limit $\Gamma\rightarrow 0$, these conductivities could be linked to the Berry curvature Oka-PRB ; Mikami . This is similar to the AHE and SHE in nondriven systems Kon-AHE ; Kon-SHE-Pt ; NA-SCD . However, for finite $\Gamma$, the anomalous Hall or spin Hall conductivity in nondriven systems contains not only the Berry-curvature term, but also the others, including the so-called Fermi-surface term Kon-AHE ; Kon-SHE-Pt ; NA-SCD ; Streda , which is distinct from the Berry-curvature term. Furthermore, the Fermi-surface term dominates the intrinsic AHE and SHE at finite $\Gamma$ Kon-AHE ; Kon-SHE-Pt ; NA-SCD ; Mizo . The similar crossover can be realized for the Hall conductivity with the magnetic field: the Hall conductivity in the strong-field case $\omega_{\textrm{c}}\tau\gg 1$ can be approximated by the Berry-curvature term TKNN , whereas that in the weak-field case $\omega_{\textrm{c}}\tau\ll 1$ is described by the Fermi-surface term Fukuyama . Here $\omega_{\textrm{c}}$ represents the energy gap between Landau levels, and $\tau$ is inversely proportional to the damping. Namely, the Berry-curvature term is dominant if the band splitting which contributes to the Hall conductivity is much larger than the broadening in the single-particle spectrum; otherwise, the Fermi-surface term is dominant. Importantly, these terms are automatically included in the conductivities derived from the Kubo formula without any additional approximation. Moreover, the limit $\Gamma\rightarrow 0$ is unrealistic in periodically driven systems because the finite $\Gamma$ is required to realize a nonequilibrium steady state with the heating due to the pump field. Therefore, we will study the time-averaged spin and charge off-diagonal dc conductivities without simplification using the Berry curvature.
V Numerical results
@@
In this section, we show the time-averaged spin and charge off-diagonal dc conductivities calculated numerically for Sr2RuO4 driven by CPL, LPL, or BCPL at $\beta=2$ or $3$. In Sec. V A, we focus on the time-averaged spin off-diagonal dc conductivities. In all the cases considered, the Onsager reciprocal relations argued in Sec. II B 2 are satisfied and the main terms are given by the antisymmetric parts. In Sec. V B, we turn to the time-averaged charge off-diagonal dc conductivities. Although the Onsager reciprocal relations argued in Sec. II A 2 are satisfied, their main terms depend on the polarization of light. In the case with CPL or LPL, the main term is the antisymmetric or symmetric part, respectively. Then, in the cases with BCPL at $(\beta,\theta)=(2,\frac{\pi}{4})$, $(2,\frac{3\pi}{4})$, $(3,\frac{\pi}{4})$, $(3,\frac{\pi}{2})$, and $(3,\frac{3\pi}{4})$, the main term is the antisymmetric part in the range of $0\leq u\leq 0.4$, whereas it is the symmetric part in the range of $0.5\leq u\leq 1$. Here $u=eA_{0}$ is the dimensionless quantity. Meanwhile, in the cases with BCPL at $(\beta,\theta)=(2,0)$, $(2,\frac{\pi}{2})$, $(2,\pi)$, $(3,0)$, and $(3,\pi)$, $\sigma_{yx}^{\textrm{C}}$ is almost vanishing in the range of $0.5\leq u\leq 1$, although its main term is the antisymmetric part in the range of $0\leq u\leq 0.4$. These results are consistent with our general arguments made in Sec. II.
[…]
@@
In the actual calculations, we set $N_{x}=N_{y}=100$, $\Delta\omega^{\prime}=0.005$ eV, $\Gamma=0.01$ eV, $T_{\textrm{b}}=0.02$ eV, and $\Omega=6$ eV. This $\Omega$ is larger than the bandwidth of the nondriven system [Fig. 5(b)], which means that the light is off-resonant. We also set $n_{\textrm{max}}=1$ except the results for $n_{\textrm{max}}=0$. Note that the calculations with $n_{\textrm{max}}=0$ include the light-induced corrections due to only the zeroth-order Bessel functions, whereas the calculations with $n_{\textrm{max}}=1$ include those due to not only the zeroth-order and first-order Bessel functions, but also the higher-order Bessel functions (see Appendix B). In Appendix D, we show that the results obtained for $n_{\textrm{max}}=1$ and $2$ are qualitatively the same in the case with BCPL at $\beta=2$, $\theta=\frac{\pi}{4}$, and $\Omega=6$ eV. The similar property holds in the case with CPL NA-FloquetSHE . Therefore, setting $n_{\textrm{max}}=1$ may be reasonable in periodically driven Sr2RuO4 at $\Omega=6$ eV.
@@
The reasons for presenting the numerical results are three fold. First, the consistency between our general arguments and numerical results supports the correctness of the results obtained in them. Second, the numerical results for a specific model may help to better understand the Onsager reciprocal relations in the periodically driven systems. Third, the numerical results can determine whether the time-averaged spin and charge off-diagonal dc conductivities are dominated by the symmetric or antisymmetric parts even in the cases with BCPL. In the cases with BCPL, the dominant terms cannot be determined from the general arguments due to the lack of a simple relation between the pump field and its time-reversal counterpart, as we have explained in Secs. II A 2 and II B 2.
@@
Figure 6: The $u(=eA_{0})$ dependences of $\sigma_{yx}^{\textrm{S}}$, $\bar{\tilde{\sigma}}_{xy}^{\textrm{S}}$, and $\tilde{\sigma}_{xy}^{\textrm{S}}$ for Sr2RuO4 driven by (a) LCPL and (b) LPL at $\Omega=6$ eV. The pump fields of LCPL and LPL have been defined in Eqs. (4) and (11), respectively. For $\sigma_{yx}^{\textrm{S}}$, $\bar{\tilde{\sigma}}_{xy}^{\textrm{S}}$, and $\tilde{\sigma}_{xy}^{\textrm{S}}$, see Figs. 4(d), 4(e), and 4(f), respectively. In the case with LPL, $\bar{\tilde{\sigma}}_{xy}^{\textrm{S}}=\tilde{\sigma}_{xy}^{\textrm{S}}$.
@@
Figure 7: (a) The $u(=eA_{0})$ dependences of $\sigma_{yx}^{\textrm{S}}$ for Sr2RuO4 driven by BCPL at $\beta=2$ and $\Omega=6$ eV with $\theta=0$, $\frac{\pi}{4}$, $\frac{\pi}{2}$, $\frac{3\pi}{4}$, and $\pi$. The pump field of BCPL has been defined in Eqs. (12) and (13). For $\sigma_{yx}^{\textrm{S}}$, see Fig. 4(d). The $u$ dependences of $\sigma_{yx}^{\textrm{S}}$, $\bar{\tilde{\sigma}}_{xy}^{\textrm{S}}$, and $\tilde{\sigma}_{xy}^{\textrm{S}}$ for Sr2RuO4 driven by BCPL at $\beta=2$ and $\Omega=6$ eV with (b) $\theta=0$, (c) $\theta=\frac{\pi}{4}$, (d) $\theta=\frac{\pi}{2}$, (e) $\theta=\frac{3\pi}{4}$, and (f) $\theta=\pi$. For $\bar{\tilde{\sigma}}_{xy}^{\textrm{S}}$, and $\tilde{\sigma}_{xy}^{\textrm{S}}$, see Figs. 4(e) and 4(f), respectively.
@@
Figure 8: The $u(=eA_{0})$ dependences of $\sigma_{yx}^{\textrm{S}}$ for Sr2RuO4 driven by BCPL at (a) $\beta=2$ and (b) $\beta=3$, $\theta=0$, and $\Omega=6$ eV with $n_{\textrm{max}}=1$ and 00. Here $n_{\textrm{max}}$ is the upper limit of the summation over the Floquet indices, i.e., $\sum_{m,l,n,q=-n_{\textrm{max}}}^{n_{\textrm{max}}}$.
V.1 Time-averaged spin off-diagonal dc conductivities
@@ V.1.1 Case with CPL or LPL
V.1.2 Cases with BCPL
@@
Figure 9: (a) The $u(=eA_{0})$ dependences of $\sigma_{yx}^{\textrm{S}}$ for Sr2RuO4 driven by BCPL at $\beta=3$ and $\Omega=6$ eV with $\theta=0$, $\frac{\pi}{4}$, $\frac{\pi}{2}$, $\frac{3\pi}{4}$, and $\pi$. The pump field of BCPL has been defined in Eqs. (12) and (13). For $\sigma_{yx}^{\textrm{S}}$, see Fig. 4(d). The $u$ dependences of $\sigma_{yx}^{\textrm{S}}$, $\bar{\tilde{\sigma}}_{xy}^{\textrm{S}}$, and $\tilde{\sigma}_{xy}^{\textrm{S}}$ for Sr2RuO4 driven by BCPL at $\beta=3$ and $\Omega=6$ eV with (b) $\theta=0$, (c) $\theta=\frac{\pi}{4}$, (d) $\theta=\frac{\pi}{2}$, (e) $\theta=\frac{3\pi}{4}$, and (f) $\theta=\pi$. For $\bar{\tilde{\sigma}}_{xy}^{\textrm{S}}$ and $\tilde{\sigma}_{xy}^{\textrm{S}}$, see Figs. 4(e) and 4(f), respectively.
@@
Figure 10: The $u(=eA_{0})$ dependences of $\sigma_{yx}^{\textrm{S}}$ for Sr2RuO4 driven by BCPL at (a) $\beta=2$ and (b) $\beta=3$, $\theta=0$, and $\Omega=6$ eV with and without the SOC. Note that in the cases with and without the SOC, $\xi=0.17$ and 00 eV, respectively.
@@
Figure 11: The $u(=eA_{0})$ dependences of $\sigma_{yx}^{\textrm{C}}$, $\bar{\sigma}_{xy}^{\textrm{C}}$, and $\sigma_{xy}^{\textrm{C}}$ for Sr2RuO4 driven by (a) LCPL and (b) LPL at $\Omega=6$ eV. The pump fields of LCPL and LPL have been defined in Eqs. (4) and (11), respectively. For $\sigma_{yx}^{\textrm{C}}$, $\bar{\sigma}_{xy}^{\textrm{C}}$, and $\sigma_{xy}^{\textrm{C}}$, see Figs. 4(a), 4(b), and 4(c), respectively. In the case with LPL, $\bar{\sigma}_{xy}^{\textrm{C}}=\sigma_{xy}^{\textrm{C}}$.
@@
Figure 12: (a) The $u(=eA_{0})$ dependences of $\sigma_{yx}^{\textrm{C}}$ for Sr2RuO4 driven by BCPL at $\beta=2$ and $\Omega=6$ eV with $\theta=0$, $\frac{\pi}{4}$, $\frac{\pi}{2}$, $\frac{3\pi}{4}$, and $\pi$. The pump field of BCPL has been defined in Eqs. (12) and (13). For $\sigma_{yx}^{\textrm{C}}$, see Fig. 4(a). The $u$ dependences of $\sigma_{yx}^{\textrm{C}}$, $\bar{\sigma}_{xy}^{\textrm{C}}$, and $\sigma_{xy}^{\textrm{C}}$ for Sr2RuO4 driven by BCPL at $\beta=2$ and $\Omega=6$ eV with (b) $\theta=0$, (c) $\theta=\frac{\pi}{4}$, (d) $\theta=\frac{\pi}{2}$, (e) $\theta=\frac{3\pi}{4}$, and (f) $\theta=\pi$. For $\bar{\sigma}_{xy}^{\textrm{C}}$ and $\sigma_{xy}^{\textrm{C}}$, see Figs. 4(b) and 4(c), respectively.
@@
Figure 13: The $u(=eA_{0})$ dependences of $\sigma_{yx}^{\textrm{C}}$ for Sr2RuO4 driven by BCPL at (a) $\beta=2$ and (b) $\beta=3$, $\theta=0$, and $\Omega=6$ eV with $n_{\textrm{max}}=1$ and 00. Here $n_{\textrm{max}}$ is the upper limit of the summation over the Floquet indices (i.e., $\sum_{m,l,n,q=-n_{\textrm{max}}}^{n_{\textrm{max}}}$).
@@
Figure 14: The $u(=eA_{0})$ dependences of $\sigma_{yx}^{\textrm{C}}$ for Sr2RuO4 driven by BCPL at $\beta=2$ and $\Omega=6$ eV with (a) $\theta=\frac{\pi}{4}$ and $\frac{3\pi}{4}$, (b) $\theta=\frac{\pi}{4}$ and $-\frac{\pi}{4}$, and (c) $\theta=\frac{3\pi}{4}$ and $\frac{5\pi}{4}$. The trajectories of the pump fields of BCPL per period $T_{\textrm{p}}$ at $\beta=2$ with (d) $\theta=\frac{\pi}{4}$ and $\frac{3\pi}{4}$, (e) $\theta=\frac{\pi}{4}$ and $-\frac{\pi}{4}$, and (f) $\theta=\frac{3\pi}{4}$ and $\frac{5\pi}{4}$. The two trajectories in panel (d) are connected by the mirror operation about the $A_{x}=0$ plane, whereas those in panel (e) or (f) are connected by the mirror operation about the $A_{y}=0$ plane.
@@
Figure 15: (a) The $u(=eA_{0})$ dependences of $\sigma_{yx}^{\textrm{C}}$ for Sr2RuO4 driven by BCPL at $\beta=3$ and $\Omega=6$ eV with $\theta=0$, $\frac{\pi}{4}$, $\frac{\pi}{2}$, $\frac{3\pi}{4}$, and $\pi$. $A_{0}$, $\beta$, and $\theta$ of BCPL have been defined in Eqs. (12) and (13). For $\sigma_{yx}^{\textrm{C}}$, see Fig. 4(a). The $u$ dependences of $\sigma_{yx}^{\textrm{C}}$, $\bar{\sigma}_{xy}^{\textrm{C}}$, and $\sigma_{xy}^{\textrm{C}}$ for Sr2RuO4 driven by BCPL at $\beta=3$ and $\Omega=6$ eV with (b) $\theta=0$, (c) $\theta=\frac{\pi}{4}$, (d) $\theta=\frac{\pi}{2}$, (e) $\theta=\frac{3\pi}{4}$, and (f) $\theta=\pi$. For $\bar{\sigma}_{xy}^{\textrm{C}}$ and $\sigma_{xy}^{\textrm{C}}$, see Figs. 4(b) and 4(c), respectively.
@@
Figure 16: The $u(=eA_{0})$ dependences of $\sigma_{yx}^{\textrm{C}}$ for Sr2RuO4 driven by BCPL at $\beta=3$ and $\Omega=6$ eV with (a) $\theta=\frac{\pi}{2}$ and $\frac{3\pi}{2}$, (b) $\theta=\frac{\pi}{4}$ and $\frac{7\pi}{4}$, and (c) $\theta=\frac{3\pi}{4}$ and $\frac{5\pi}{4}$. The trajectories of the pump fields of BCPL per period $T_{\textrm{p}}$ at $\beta=3$ with (d) $\theta=\frac{\pi}{2}$ and $\frac{3\pi}{2}$, (e) $\theta=\frac{\pi}{4}$ and $\frac{7\pi}{4}$, and (f) $\theta=\frac{3\pi}{4}$ and $\frac{5\pi}{4}$. The two trajectories in each panel are connected by the mirror operation about the $A_{x}=0$ or $A_{y}=0$ plane.
@@
Figure 17: The $u(=eA_{0})$ dependences of $\sigma_{yx}^{\textrm{C}}$ for Sr2RuO4 driven by BCPL at (a) $\beta=2$ and (b) $\beta=3$ and $\Omega=6$ eV with $\theta=0$, $\frac{\pi}{4}$, $\frac{\pi}{2}$, $\frac{3\pi}{4}$, and $\pi$ in the absence of SOC.

### v2

@@ 4 Floquet linear-response theory
and performing the similar procedures to those used to determine $G^{r}_{a\sigma b\sigma^{\prime}}(\mathbf{k},\omega)]_{mn}$’s. Namely, the group velocities and Green’s functions appearing in Eqs. (89) and (90) are calculated with $\mathbf{A}_{\textrm{pump}}(-t)$ in the similar way to those with $\mathbf{A}_{\textrm{pump}}(t)$. For $\bar{\epsilon}_{ab}(\mathbf{k})]_{mn}$ of Sr2RuO4 driven by BCPL, see Appendix C.
4.3 General remarks about the applicability
@@
Then, our theory has wider applicability than the theories using a high-frequency expansion. This expansion has been often used to analyze many periodically driven systems [3, 58, 59, 8, 30]. It may be sufficient for a periodically driven electron system if the light frequency is much larger than the bandwidth of the system. Meanwhile, our theory does not have such a restriction because the Floquet theory used here is free from the constraint on the frequency.
@@
The applicability of our theory is also wider than that of the theories in which the time-averaged off-diagonal dc conductivities are expressed in terms of the Berry curvature. In the limit $\Gamma\rightarrow 0$, these conductivities could be linked to the Berry curvature [7, 8]. This is similar to the AHE and SHE in nondriven systems [60, 61, 62]. However, for finite $\Gamma$, the anomalous Hall or spin Hall conductivity in nondriven systems contains not only the Berry-curvature term, but also the others, including the so-called Fermi-surface term [60, 61, 62, 63], which is distinct from the Berry-curvature term. Furthermore, the Fermi-surface term dominates the intrinsic AHE and SHE at finite $\Gamma$ [60, 61, 62, 46]. The similar crossover can be realized for the Hall conductivity with the magnetic field: the Hall conductivity in the strong-field case $\omega_{\textrm{c}}\tau\gg 1$ can be approximated by the Berry-curvature term [64], whereas that in the weak-field case $\omega_{\textrm{c}}\tau\ll 1$ is described by the Fermi-surface term [65]. Here $\omega_{\textrm{c}}$ represents the energy gap between Landau levels, and $\tau$ is inversely proportional to the damping. Namely, the Berry-curvature term is dominant if the band splitting which contributes to the Hall conductivity is much larger than the broadening in the single-particle spectrum; otherwise, the Fermi-surface term is dominant. Importantly, these terms are automatically included in the conductivities derived from the Kubo formula without any additional approximation. Moreover, the limit $\Gamma\rightarrow 0$ is unrealistic in periodically driven systems because the finite $\Gamma$ is required to realize a nonequilibrium steady state with the heating due to the pump field. Therefore, we will study the time-averaged spin and charge off-diagonal dc conductivities without simplification using the Berry curvature.
5 Numerical results
@@
In this section, we show the time-averaged spin and charge off-diagonal dc conductivities calculated numerically for Sr2RuO4 driven by CPL, LPL, or BCPL at $\beta=2$ or $3$. In Sect. 5.1, we focus on the time-averaged spin off-diagonal dc conductivities. In all the cases considered, the Onsager reciprocal relations argued in Sect. 2.2.2 are satisfied and the main terms are given by the antisymmetric parts. In Sect. 5.2, we turn to the time-averaged charge off-diagonal dc conductivities. Although the Onsager reciprocal relations argued in Sect. 2.1.2 are satisfied, their main terms depend on the polarization of light. In the case with CPL or LPL, the main term is the antisymmetric or symmetric part, respectively. Then, in the cases with BCPL at $(\beta,\theta)=(2,\frac{\pi}{4})$, $(2,\frac{3\pi}{4})$, $(3,\frac{\pi}{4})$, $(3,\frac{\pi}{2})$, and $(3,\frac{3\pi}{4})$, the main term is the antisymmetric part in the range of $0\leq u\leq 0.4$, whereas it is the symmetric part in the range of $0.5\leq u\leq 1$. Here $u=eA_{0}$ is the dimensionless quantity. Meanwhile, in the cases with BCPL at $(\beta,\theta)=(2,0)$, $(2,\frac{\pi}{2})$, $(2,\pi)$, $(3,0)$, and $(3,\pi)$, $\sigma_{yx}^{\textrm{C}}$ is almost vanishing in the range of $0.5\leq u\leq 1$, although its main term is the antisymmetric part in the range of $0\leq u\leq 0.4$. These results are consistent with our general arguments made in Sect. 2.
We numerically evaluated Eqs. (57), (58), and (87)–(90) using the following four procedures. First, we calculated the momentum summation by dividing the first Brillouin zone into a $N_{x}\times N_{y}$ mesh. Second, we performed the frequency integration using $\int_{-\Omega/2}^{\Omega/2}d\omega^{\prime}F(\omega^{\prime})\approx\sum_{s=0}^{W-1}\Delta\omega^{\prime}F(\omega^{\prime}_{s})$, where $\omega^{\prime}_{s}=-\Omega/2+s\Delta\omega^{\prime}$, and $\omega^{\prime}_{W}=\Omega/2$. Third, we calculated the frequency derivatives of the Green’s functions by using $\frac{\partial F(\omega^{\prime})}{\partial\omega^{\prime}}\approx\frac{F(\omega^{\prime}+\Delta\omega^{\prime})-F(\omega^{\prime}-\Delta\omega^{\prime})}{2\Delta\omega^{\prime}}$. Fourth, we replaced the summations over the Floquet indices, $\sum_{m,l,n,q=-\infty}^{\infty}$, by $\sum_{m,l,n,q=-n_{\textrm{max}}}^{n_{\textrm{max}}}$.
@@
In the actual calculations, we set $N_{x}=N_{y}=100$, $\Delta\omega^{\prime}=0.005$ eV, $\Gamma=0.01$ eV, $T_{\textrm{b}}=0.02$ eV, and $\Omega=6$ eV. This $\Omega$ is larger than the bandwidth of the nondriven system [Fig. 5(b)], which means that the light is off-resonant. We also set $n_{\textrm{max}}=1$ except the results for $n_{\textrm{max}}=0$. Note that the calculations with $n_{\textrm{max}}=0$ include the light-induced corrections due to only the zeroth-order Bessel functions, whereas the calculations with $n_{\textrm{max}}=1$ include those due to not only the zeroth-order and first-order Bessel functions, but also the higher-order Bessel functions (see Appendix B). In Appendix D, we show that the results obtained for $n_{\textrm{max}}=1$ and $2$ are qualitatively the same in the case with BCPL at $\beta=2$, $\theta=\frac{\pi}{4}$, and $\Omega=6$ eV. The similar property holds in the case with CPL [11]. Therefore, setting $n_{\textrm{max}}=1$ may be reasonable in periodically driven Sr2RuO4 at $\Omega=6$ eV.
@@
The reasons for presenting the numerical results are three fold. First, the consistency between our general arguments and numerical results supports the correctness of the results obtained in them. Second, the numerical results for a specific model may help to better understand the Onsager reciprocal relations in the periodically driven systems. Third, the numerical results can determine whether the time-averaged spin and charge off-diagonal dc conductivities are dominated by the symmetric or antisymmetric parts even in the cases with BCPL. In the cases with BCPL, the dominant terms cannot be determined from the general arguments due to the lack of a simple relation between the pump field and its time-reversal counterpart, as we have explained in Sects. 2.1.2 and 2.2.2.
@@
Figure 6: (Color online) The $u(=eA_{0})$ dependences of $\sigma_{yx}^{\textrm{S}}$, $\bar{\tilde{\sigma}}_{xy}^{\textrm{S}}$, and $\tilde{\sigma}_{xy}^{\textrm{S}}$ for Sr2RuO4 driven by (a) LCPL and (b) LPL at $\Omega=6$ eV. The pump fields of LCPL and LPL have been defined in Eqs. (4) and (11), respectively. For $\sigma_{yx}^{\textrm{S}}$, $\bar{\tilde{\sigma}}_{xy}^{\textrm{S}}$, and $\tilde{\sigma}_{xy}^{\textrm{S}}$, see Figs. 4(d), 4(e), and 4(f), respectively. In the case with LPL, $\bar{\tilde{\sigma}}_{xy}^{\textrm{S}}=\tilde{\sigma}_{xy}^{\textrm{S}}$.
@@
Figure 7: (Color online) (a) The $u(=eA_{0})$ dependences of $\sigma_{yx}^{\textrm{S}}$ for Sr2RuO4 driven by BCPL at $\beta=2$ and $\Omega=6$ eV with $\theta=0$, $\frac{\pi}{4}$, $\frac{\pi}{2}$, $\frac{3\pi}{4}$, and $\pi$. The pump field of BCPL has been defined in Eqs. (12) and (13). For $\sigma_{yx}^{\textrm{S}}$, see Fig. 4(d). The $u$ dependences of $\sigma_{yx}^{\textrm{S}}$, $\bar{\tilde{\sigma}}_{xy}^{\textrm{S}}$, and $\tilde{\sigma}_{xy}^{\textrm{S}}$ for Sr2RuO4 driven by BCPL at $\beta=2$ and $\Omega=6$ eV with (b) $\theta=0$, (c) $\theta=\frac{\pi}{4}$, (d) $\theta=\frac{\pi}{2}$, (e) $\theta=\frac{3\pi}{4}$, and (f) $\theta=\pi$. For $\bar{\tilde{\sigma}}_{xy}^{\textrm{S}}$, and $\tilde{\sigma}_{xy}^{\textrm{S}}$, see Figs. 4(e) and 4(f), respectively.
@@
Figure 8: (Color online) The $u(=eA_{0})$ dependences of $\sigma_{yx}^{\textrm{S}}$ for Sr2RuO4 driven by BCPL at (a) $\beta=2$ and (b) $\beta=3$, $\theta=0$, and $\Omega=6$ eV with $n_{\textrm{max}}=1$ and 0. Here $n_{\textrm{max}}$ is the upper limit of the summation over the Floquet indices, i.e., $\sum_{m,l,n,q=-n_{\textrm{max}}}^{n_{\textrm{max}}}$.
5.1 Time-averaged spin off-diagonal dc conductivities
@@ 5.1.1 Case with CPL or LPL
5.1.2 Cases with BCPL
@@
Figure 9: (Color online) (a) The $u(=eA_{0})$ dependences of $\sigma_{yx}^{\textrm{S}}$ for Sr2RuO4 driven by BCPL at $\beta=3$ and $\Omega=6$ eV with $\theta=0$, $\frac{\pi}{4}$, $\frac{\pi}{2}$, $\frac{3\pi}{4}$, and $\pi$. The pump field of BCPL has been defined in Eqs. (12) and (13). For $\sigma_{yx}^{\textrm{S}}$, see Fig. 4(d). The $u$ dependences of $\sigma_{yx}^{\textrm{S}}$, $\bar{\tilde{\sigma}}_{xy}^{\textrm{S}}$, and $\tilde{\sigma}_{xy}^{\textrm{S}}$ for Sr2RuO4 driven by BCPL at $\beta=3$ and $\Omega=6$ eV with (b) $\theta=0$, (c) $\theta=\frac{\pi}{4}$, (d) $\theta=\frac{\pi}{2}$, (e) $\theta=\frac{3\pi}{4}$, and (f) $\theta=\pi$. For $\bar{\tilde{\sigma}}_{xy}^{\textrm{S}}$ and $\tilde{\sigma}_{xy}^{\textrm{S}}$, see Figs. 4(e) and 4(f), respectively.
@@
Figure 10: (Color online) The $u(=eA_{0})$ dependences of $\sigma_{yx}^{\textrm{S}}$ for Sr2RuO4 driven by BCPL at (a) $\beta=2$ and (b) $\beta=3$, $\theta=0$, and $\Omega=6$ eV with and without the SOC. Note that in the cases with and without the SOC, $\xi=0.17$ and 0 eV, respectively.
@@
Figure 11: (Color online) The $u(=eA_{0})$ dependences of $\sigma_{yx}^{\textrm{C}}$, $\bar{\sigma}_{xy}^{\textrm{C}}$, and $\sigma_{xy}^{\textrm{C}}$ for Sr2RuO4 driven by (a) LCPL and (b) LPL at $\Omega=6$ eV. The pump fields of LCPL and LPL have been defined in Eqs. (4) and (11), respectively. For $\sigma_{yx}^{\textrm{C}}$, $\bar{\sigma}_{xy}^{\textrm{C}}$, and $\sigma_{xy}^{\textrm{C}}$, see Figs. 4(a), 4(b), and 4(c), respectively. In the case with LPL, $\bar{\sigma}_{xy}^{\textrm{C}}=\sigma_{xy}^{\textrm{C}}$.
@@
Figure 12: (Color online) (a) The $u(=eA_{0})$ dependences of $\sigma_{yx}^{\textrm{C}}$ for Sr2RuO4 driven by BCPL at $\beta=2$ and $\Omega=6$ eV with $\theta=0$, $\frac{\pi}{4}$, $\frac{\pi}{2}$, $\frac{3\pi}{4}$, and $\pi$. The pump field of BCPL has been defined in Eqs. (12) and (13). For $\sigma_{yx}^{\textrm{C}}$, see Fig. 4(a). The $u$ dependences of $\sigma_{yx}^{\textrm{C}}$, $\bar{\sigma}_{xy}^{\textrm{C}}$, and $\sigma_{xy}^{\textrm{C}}$ for Sr2RuO4 driven by BCPL at $\beta=2$ and $\Omega=6$ eV with (b) $\theta=0$, (c) $\theta=\frac{\pi}{4}$, (d) $\theta=\frac{\pi}{2}$, (e) $\theta=\frac{3\pi}{4}$, and (f) $\theta=\pi$. For $\bar{\sigma}_{xy}^{\textrm{C}}$ and $\sigma_{xy}^{\textrm{C}}$, see Figs. 4(b) and 4(c), respectively.
@@
Figure 13: (Color online) The $u(=eA_{0})$ dependences of $\sigma_{yx}^{\textrm{C}}$ for Sr2RuO4 driven by BCPL at (a) $\beta=2$ and (b) $\beta=3$, $\theta=0$, and $\Omega=6$ eV with $n_{\textrm{max}}=1$ and 0. Here $n_{\textrm{max}}$ is the upper limit of the summation over the Floquet indices (i.e., $\sum_{m,l,n,q=-n_{\textrm{max}}}^{n_{\textrm{max}}}$).
@@
Figure 14: (Color online) The $u(=eA_{0})$ dependences of $\sigma_{yx}^{\textrm{C}}$ for Sr2RuO4 driven by BCPL at $\beta=2$ and $\Omega=6$ eV with (a) $\theta=\frac{\pi}{4}$ and $\frac{3\pi}{4}$, (b) $\theta=\frac{\pi}{4}$ and $\frac{\pi}{4}$, and (c) $\theta=\frac{3\pi}{4}$ and $\frac{5\pi}{4}$. The trajectories of the pump fields of BCPL per period $T_{\textrm{p}}$ at $\beta=2$ with (d) $\theta=\frac{\pi}{4}$ and $\frac{3\pi}{4}$, (e) $\theta=\frac{\pi}{4}$ and $\frac{\pi}{4}$, and (f) $\theta=\frac{3\pi}{4}$ and $\frac{5\pi}{4}$. The two trajectories in panel (d) are connected by the mirror operation about the $A_{x}=0$ plane, whereas those in panel (e) or (f) are connected by the mirror operation about the $A_{y}=0$ plane.
@@
Figure 15: (Color online) (a) The $u(=eA_{0})$ dependences of $\sigma_{yx}^{\textrm{C}}$ for Sr2RuO4 driven by BCPL at $\beta=3$ and $\Omega=6$ eV with $\theta=0$, $\frac{\pi}{4}$, $\frac{\pi}{2}$, $\frac{3\pi}{4}$, and $\pi$. $A_{0}$, $\beta$, and $\theta$ of BCPL have been defined in Eqs. (12) and (13). For $\sigma_{yx}^{\textrm{C}}$, see Fig. 4(a). The $u$ dependences of $\sigma_{yx}^{\textrm{C}}$, $\bar{\sigma}_{xy}^{\textrm{C}}$, and $\sigma_{xy}^{\textrm{C}}$ for Sr2RuO4 driven by BCPL at $\beta=3$ and $\Omega=6$ eV with (b) $\theta=0$, (c) $\theta=\frac{\pi}{4}$, (d) $\theta=\frac{\pi}{2}$, (e) $\theta=\frac{3\pi}{4}$, and (f) $\theta=\pi$. For $\bar{\sigma}_{xy}^{\textrm{C}}$ and $\sigma_{xy}^{\textrm{C}}$, see Figs. 4(b) and 4(c), respectively.
@@
Figure 16: (Color online) The $u(=eA_{0})$ dependences of $\sigma_{yx}^{\textrm{C}}$ for Sr2RuO4 driven by BCPL at $\beta=3$ and $\Omega=6$ eV with (a) $\theta=\frac{\pi}{2}$ and $\frac{3\pi}{2}$, (b) $\theta=\frac{\pi}{4}$ and $\frac{7\pi}{4}$, and (c) $\theta=\frac{3\pi}{4}$ and $\frac{5\pi}{4}$. The trajectories of the pump fields of BCPL per period $T_{\textrm{p}}$ at $\beta=3$ with (d) $\theta=\frac{\pi}{2}$ and $\frac{3\pi}{2}$, (e) $\theta=\frac{\pi}{4}$ and $\frac{7\pi}{4}$, and (f) $\theta=\frac{3\pi}{4}$ and $\frac{5\pi}{4}$. The two trajectories in each panel are connected by the mirror operation about the $A_{x}=0$ or $A_{y}=0$ plane.
@@
Figure 17: (Color online) The $u(=eA_{0})$ dependences of $\sigma_{yx}^{\textrm{C}}$ for Sr2RuO4 driven by BCPL at (a) $\beta=2$ and (b) $\beta=3$ and $\Omega=6$ eV with $\theta=0$, $\frac{\pi}{4}$, $\frac{\pi}{2}$, $\frac{3\pi}{4}$, and $\pi$ in the absence of SOC.